Show calculations setup please.

A sample of methan gas, CH4, occupies 4.50L at a temperature of 20.0 degrees C. If the pressure is held constant, what will be the volume of the gas at 100. degrees C?
Answer #1

Given:

Vi = 4.50 L

Ti= 20.0 oC = (20.0+273) K = 293 K

Tf= 100.0 oC = (100.0+273) K = 373.0 K

use:

Vi/Ti = Vf/Tf

4.50 L / 293 K = Vf / 373 K

Vf = 5.73 L

Answer: 5.73 L